India is the second most populous country in the world, with a population of over 1 billion people. Although the government has offered various incentives for population control, some argue that the birth rate, especially in rural India, is still too high to be sustainable. A demographer assumes the following probability distribution of the household size in India 0.02 0.09 0.14 0.21 0.29 0.16 0.07 0.02 a. What is the probability that there are less than 5 members in a household in India? (Round your answer to 2 decimal places.) Probability b. What is the probability that there are 5 or more members in a household in india? (Round your answer to 2 decimal places.) Probability c. What is the probability that the number of members in a household in India is strictly between 3 and 6? Round your answer to 2 decimal places.) Probability References eBook & Resources Worksheet Type here to search
Answer #1

a) P ( X < 5) = P (X=1) + P(X=2)+P(X=3) + P(X=4)

= 0.02+ 0.09+0.14+0.21

=0.46

b) P(X>=5) = P(X=5) + P(X=6) +....P(X=8)

= 0.29+0.16+0.07+0.02

= 0.54

c) P(3<= X <=6) = 0.14+ 0.21+0.29+ 0.16 = 0.80